OpenStudy (anonymous):

solve the following quadratic equations by factoring. set equations to "0" first tho. show work please 3x^2+16x=35

OpenStudy (anonymous):

3x^2 + 16x - 35 = 0 x = -7, 5/3 by the quadratic formula

OpenStudy (anonymous):

i cant use the quadratic formula on this problem only by factoring.

OpenStudy (anonymous):

(x+7) (3 x-5) = 0 x+7 = 0 3x-5 = 0 x = -7, 5/3

OpenStudy (anonymous):

thanks